# A Mixed-Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling

### v3

@@ 4. Methodology
We adapted a sequential, exploratory mixed-methods design (Creswell et al., 2003) consisting of three phases. We describe the first phase in Section 0, where we used a screening survey to recruit developers for semi-structured interviews. In the second phase, after interviewing each eligible candidate, the first three authors conducted a thematic analysis (Miles et al., 2020) of the interview transcripts. We describe this process in Section 0. We describe the third phase in Section 0,where we used the results of our qualitative analysis to create a community survey. We chose this combination of methods because it allowed us to evaluate the generalizability of prior qualitative findings and to provide a missing qualitative perspective on interoperation and tool use.
@@
(c) Interview participants’ affiliations ("Affil."), their years of experience using Rust ("Years"), and the number of coding decisions made for their interview (Decs.)
ID Affil. Years Decs.
1 I 2 56
2 I,R 10 76
3 I,O 5 93
4 I,O 8 71
5 I,O 2 61
6 A 3 71
7 I 6 57
8 I 5 49
9 O 2 97
10 O 2 59
11 I 5 86
12 I,A 6 121
13 I 10 92
14 R,O,I 7 94
15 I 1.5 58
16 I 3 20
17 A 4 64
18 I 3 82
19 I 4 13
I = Industry  R = Rust Team
A = Academia  O = Open Source
(d) Codes for the theme "Why unsafe?", which describes participants’ motivations for using unsafe code.
@@ 19 I 4 13
(e) A survey question with a direct correspondence to the codes shown in Figure 0.
@@ 4.1. Interviews
We used participants’ responses to the screening survey to guide each interview. We provide a sample of our interview protocol in Figure 0. Certain questions were gated based on prior responses. If an candidate answered "No" to questions #0 and #0 of Figure 0, then they did not use Rust in multi-language applications, so we did not ask them questions #0-0 of the screening survey or question #0 of the interview protocol, since each of these questions involve foreign functions.
The first author conducted all 19 interviews remotely over Zoom between May and July of 2023. We used snowball sampling (Patton, 1990) to expand our pool of participants beyond the group that responded to our initial calls to participate. If a participant shared new information that we had not yet learned, then we prompted them to invite any acquaintances with relevant experience to complete our screening survey. We stopped recruiting new participants after we had reached “data saturation” as defined by Saunders et al. (Saunders et al., 2018); we began to hear the same or similar responses to each of our questions. We transcribed the audio recordings of the interviews using Whisper (Radford et al., 2023), and the first author reviewed every transcript to correct errors and label whether utterances came from the interviewer or the interviewee.
@@ 4.2. Qualitative Analysis
We conducted an inductive, thematic analysis of our interview transcripts following best practices by Miles et al. (Miles et al., 2020). To begin, the first three authors conducted open coding of random samples of quotes taken from each interview. To create a sample, the first author picked random character indices within each of the interview transcripts and chose a snippet of one of the interviewees’ responses near each index. If all responses at that index had been sampled before, then the first author generated a new random index and repeated this process until a new quote was found. We coded three samples of quotes, and we met after each sample to resolve disagreements and record new codes within a shared codebook.
Next, the first and second authors used a card-sorting (Braun and Clarke, 2006) exercise to identify higher-level themes among the codes that we had generated in the first phase. We created cards for each of the quotes randomly sampled in the first phase, as well as for a set of quotes that were highlighted by the first author when reviewing each transcript. We used the themes that we created from this process to organize our codebook, and we removed redundant or irrelevant codes. Then, the first and second authors used this codebook to conduct closed coding of 7 additional random samples of quotes, meeting after coding each sample to resolve disagreements and update the codebook. We used the same process as before to create samples, but instead of treating each interview transcript separately, we sampled quotes from a single document containing the concatenated text of every interview transcript. Table 0 shows one of the themes that we created during this process, as well as its constituent codes.
After this refinement stage, the first and second authors divided the set of transcripts between each other to code in full, independently, using ATLAS.ti Web (GmbH., 2024). The first author coded 11 transcripts, while the second author coded the remaining 8 transcripts. Throughout this process, both authors continued to update a shared codebook with additional codes and themes to describe new observations. After all transcripts had been coded by one of the two authors, the first author conducted an additional coding pass for every transcript and resolved each of their disagreements with the second author. The fourth column of Table 0 shows the final number of coding decisions made for each participants’ transcript. The complete codebook and all coding decisions are included in our replication package.
@@ 4.3. Survey
We constructed the community survey after we had completed coding every interview, but before the first author had finished auditing coding decisions. Each of the first three authors independently assembled a list of survey questions, which the first author assembled into a single survey in Qualtrics (Qualtrics, 2024). Excluding consent and eligibility, the survey had 85 yes/no, multiple-choice, and Likert scale questions, which were grouped into 26 sections. To avoid priming participants to select certain responses, we randomly shuffled the order of multiple choice and yes/no options and randomly reversed the order of Likert scale items. Similar to our screening survey, certain questions were gated based on prior responses. We provide a complete copy of our survey, including its branching logic, in Section 4 of our Appendix.
@@
Questions were either descriptive or exploratory. Figure 0 shows an example of a descriptive question. Each of its response options correspond to the codes shown in Figure 0 for the theme “Why unsafe?”. Exploratory questions arose after we had finished conducting interviews, and they did not have a direct correspondence to the themes that we identified during qualitative analysis. For example, Höltervennhoff et al. (Höltervennhoff et al., 2023)’s study was published concurrent to our investigation, and three of their participants mentioned using either cargo-audit or cargo-deny to detect vulnerabilities in their dependencies. None of our interview participants had mentioned using these tools, but we included a survey question asking whether respondents had used them before.
We also relaxed our eligibility criteria for the survey to include individuals who would have been eligible for Höltervenhoff et al. (Höltervennhoff et al., 2023)’s study. Candidates were eligible to participate in their study if they had committed unsafe code to GitHub, but there was no indication if they had done so regularly. Our interview participants needed to have “regularly” engaged with unsafe Rust, but our survey respondents were eligible if they had “written, edited, read, audited, or engaged with unsafe Rust code in any way.”
@@
Two of our interview participants piloted an initial draft of the survey, and we incorporated their feedback into the final version. We distributed our community survey on September 22nd, 2023 to the same acquaintances and communities where we distributed our screening survey. We also published a link to our survey in the September 28th edition of the “This Week in Rust” newsletter, and we advertised the survey during a presentation and poster sessions at an industry partners conference hosted by our institution. At the end of the survey, participants could provide a link to their profile on either GitHub or the Rust Programming Language Forum to have a chance to receive one of two \$250 gift cards. Only participants with public account activity prior to the start date of the survey were eligible for compensation. We closed the survey on October 9th, 2023.
@@ 5. Ethics & Threats to Validity
All procedures in this study were approved by our institution’s IRB. Prior to each intervention, participants were read or presented a script outlining the procedures of the study and they were asked to affirm consent. Participants were reminded to avoid sharing personally identifying information in their responses, and interview participants were asked to remain in a private space. The first author reviewed the contents of the replication package and redacted all direct (e.g. name, affiliation) and indirect (e.g. position, project) personal identifiers.
@@
Recruitment via snowball sampling and through online communities may bias our results toward certain areas of the Rust community. We attempted to maintain a diverse survey and interview population by sampling from multiple forums. However, both Reddit and “This Week in Rust” were over-represented compared to the other communities where we distributed our survey. We cannot determine if our interview population has a similar bias, since we did not record where interview participants had first learned about our study. Additionally, the gift cards that we offered for compensation are most easily redeemable in North America, which may have dissuaded certain individuals from participating in either stage of our study. We did not collect location data from participants, so we cannot determine if our results are biased toward individuals from any particular area.
Respondents were eligible to complete our survey if they had one year of experience and had “engaged” with unsafe Rust in any way. A subset may have been eligible without having any significant practical experience with unsafe code. We mitigated this by gating questions based on whether participants had used certain features. When relevant, we compare results from the entire population with results from the subset who “regularly” wrote or edited unsafe code, which matches the eligibility criteria for our interviews. This criteria may have introduced self-selection bias among our interview participants, since individual who frequently use unsafe code may be more likely to perceive that it is necessary for their purposes.
@@
The validity of our results depends on participants’ ability to accurately self-report their behavior, and what developers perceive about their use of unsafe code may differ from how it is used in practice. Our qualitative results are unavoidably influenced by the individual biases of each investigator. We attempted to mitigate this through auditing coding decisions and meeting frequently to resolve disagreements. When we distributed our community survey, we configured Qualtrics (Qualtrics, 2024) to reject multiple responses from the same individual using IP information (which was not recorded or visible to investigators), RelevantID (RelevantID, 2024) and reCaptcha (Liu, 2018). We did not enable these features for the initial screening survey. The overwhelming majority of interview participants enabled their camera during interviews, but a few did not; they were not prompted or required to do so. We rely on the baseline assumption that every participant was honest about their identity, but we could not independently verify this in every situation.
@@ 6. Results
We found that participants used unsafe code for ergonomics, performance, and when they perceived that there was no safe alternative. They attempted to keep unsafe code minimal and encapsulated, but many participants were still uncertain about its correctness. Most participants had called foreign functions, and non-trivial differences between Rust and other languages made these calls difficult to encapsulate in certain situations. Miri was the most popular bug finding tool. However, while more than 60% of respondents had used Miri before, its performance and lack of support for key features deterred 69% of these respondents from using it again. Our participants would benefit from tools that can provide guarantees of soundness and that have robust support for multi-language applications.
We report the demographics of interview and survey participants in Section 0. Then, we describe how these developers reason about foreign functions in Section 0 (RQ1), and we discuss the tooling they use in Section 0 (RQ2). We describe developers’ motivations for using unsafe code in Section 0 (RQ3), and we report how they reason about encapsulation in Section 0 (RQ4). We summarize key results and provide suggestions for tool designers at the end of each section.
@@ 6.1. Demographics
[…]
Years of Experience
Location # SE C C++ Rust
Rust Forums 16 17.3 ± 10.1 7.5 ± 6.8 5.6 ± 6.2 4.5 ± 2.4
Rust Discord 10 8 ± 3.7 2.7 ± 2.8 3.4 ± 2.3 3.6 ± 2
Reddit (/r/rust) 83 13.9 ± 9.1 7.6 ± 7.2 6 ± 6.3 4.1 ± 2.2
"This Week in Rust" 52 15.1 ± 9.7 6.6 ± 7.7 5.7 ± 6.9 4.1 ± 2.3
Other 3 23.7 ± 15.8 17.3 ± 14.6 9 ± 5.6 3.7 ± 1.5
All 160 14.5 ± 9.5 7.2 ± 7.5 5.8 ± 6.3 4.1 ± 2.2
@@
We invited all 42 eligible candidates to participate in interviews, and we interviewed 19 who responded to our invitation. Table 0 summarizes interview participants’ demographic information. There were 15 who had experience in industry, 6 in open source development, 3 in academia, and 2 were members of the Rust Team. They had 4.7 years of experience with Rust on average, and the most experienced participant had been using Rust for 10 years. We did not collected any additional demographic information from our interview participants.
@@
We received 370 survey responses, of which 242 were complete, 203 met our eligibility criteria, and 160 passed each of Qualtrics’ (Qualtrics, 2024) measures for fraud detection. Hereafter, when we refer to “respondents,” we include only the 160 that passed every one of our checks. We collected several categories of demographic information from survey respondents. The majority were young, educated, identified as male, and had experience in either industry or open-source development. Nearly 66% had completed at least a bachelor’s degree, and 30% had completed a graduate degree. Most respondents (83%) were younger than 40 years old, and 56% were between the ages of 18 and 29. We followed Spiel et al.’s (Spiel et al., 2019) guidelines for ethically surveying respondents’ gender; 80% identified as “Man”, 6% identified as “Non-binary”, 3% identified as “Woman” and 9% did not disclose their gender. Table 0 summarizes respondents’ years of experience with the field of software engineering, Rust, and other systems languages. On average, respondents had more than a decade of experience in software engineering, more than 5 years of experience in both C and C++, and just over 4 years of experience using Rust.
@@
Table 3. How frequently survey respondents “engaged with” and wrote unsafe code. Percentages on the left include “Less than once a year” and “Yearly”, while those on the right include “Weekly” and “Daily”. The number of respondents in the sample is listed in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
[…]
[Uncaptioned image]23%41%(160)
@@
3b How frequently do you write new Rust code or edit existing Rust code within an unsafe block or function?
[Uncaptioned image]31%26%(160)
[…]
@@
Participants had a variety of roles within the Rust community and differing levels of engagement with unsafe code. Among survey respondents, 74% had committed unsafe code to a public GitHub repository, and 66% had contributed to a published crate. Our interview population had regularly written or edited unsafe code, but 47% of our survey population did not. Since “regularly” is subjective, we asked the questions shown in Table 0 to provide a more specific time-frame and to distinguish editing unsafe code from practices that do not require code contributions, such as auditing. Only 26% (6.1) of survey respondents would write or edit unsafe code more than once a month, but 41% (6.1) would engage with it at this frequency.

### v4

@@ 3. Methodology
We investigated how Rust developers engage with unsafe code by adapting a sequential, exploratory mixed methods design (Creswell et al., 2003), which consisted of three phases. We describe the first phase in Section 0, where we used a screening survey to recruit developers for semi-structured interviews. In the second phase, after interviewing each eligible candidate, the first three authors conducted a thematic analysis (Miles et al., 2020) of the interview transcripts. We describe this process in Section 0. We describe the third phase in Section 0, where we used the results of our qualitative analysis to create a community survey. We chose this combination of methods because it allowed us to evaluate the generalizability of prior qualitative findings and to provide a missing qualitative perspective on interoperation and tool use.
@@
(c) Codes for the theme "Why unsafe?", which describes participants’ motivations for using unsafe code.
@@
(d) A survey question with a direct correspondence to the codes shown in Figure 0.
@@ 3.1. Interviews
We used participants’ responses to the screening survey to guide each interview. We provide a sample of our interview protocol in Figure 0. Certain questions were gated based on prior responses. If an candidate answered "No" to questions #0 and #0 of Figure 0, then they did not use Rust in multilanguage applications, so we did not ask them questions #0-0 of the screening survey or question #0 of the interview protocol, since each of these questions involve foreign functions.
The first author conducted all 19 interviews remotely over Zoom between May and July of 2023. Participants were not required to enable their camera. We used snowball sampling (Patton, 1990) to expand our pool of participants beyond the group that responded to our initial calls to participate. If a participant shared new information that we had not learned yet, then we prompted them to invite any acquaintances with relevant experience to complete our screening survey. We stopped recruiting new participants after we had reached saturation. Our concept of saturation corresponds to Saunders’ et al (Saunders et al., 2018)’s definition of data saturation as a subjective judgement made during data collection before analysis. We knew that we had reached this point after we began to “hear the same comments again and again” (Grady, 1998) and had informally identified several themes from our interviews. Two themes were particularly notable at this stage; certain participants felt that Rust’s Box could be cumbersome due to its noalias semantics (Section 0), and that Miri was difficult to use in multilanguage applications (Section 0). Our sample size (19) is comparable to other interview studies with Rust developers (Fulton et al., 2021; Höltervennhoff et al., 2023). We transcribed the audio recordings of the interviews using Whisper (Radford et al., 2023), and the first author reviewed every transcript to correct errors and label whether the utterances came from the interviewer or the interviewee.
@@ 3.2. Qualitative Analysis
We conducted an inductive, thematic analysis of our interview transcripts following best-practices by Miles, Huberman, and Saldaña (Miles et al., 2020). To begin, the first three authors conducted open coding of random samples of quotes taken from each interview. To create a sample, the first author picked random character indices within each of the interview transcripts and chose a snippet of one of the interviewees’ responses near each index. If all responses at that index had been sampled before, then the first author generated a new random index and repeated this process until a new quote was found. We coded three samples of quotes and met after each sample to resolve disagreements and record new codes within a shared codebook.
Next, the first and second authors used a card-sorting (Braun and Clarke, 2006) exercise to identify higher-level themes among the codes that we had generated in the first phase. We created cards for each of the quotes randomly sampled in the first phase, as well as for a set of quotes that were highlighted by the first author when reviewing each transcript. We used the themes that we created from this process to organize our codebook, and we removed redundant or irrelevant codes. The first and second authors used this codebook to conduct closed coding of seven additional random samples of quotes, meeting after coding each sample to resolve disagreements and update the codebook. We used the same process as before to create samples, but instead of treating each interview transcript separately, we sampled quotes from a single document containing the concatenated text of every interview transcript. Table 0 shows one of the themes that we created during this process, as well as its constituent codes.
After this refinement stage, the first and second authors divided the set of transcripts between each other to code in full, independently, using ATLAS.ti Web (GmbH., 2024). The first author coded 11 transcripts, while the second author coded the remaining 8 transcripts. Throughout this process, both authors continued to update a shared codebook with additional codes and themes to describe new observations. After all transcripts had been coded by one of the two authors, the first author conducted an additional coding pass for every transcript and resolved each of their disagreements with the second author. The complete codebook and all coding decisions are included in our replication package.
@@ 3.3. Survey
We constructed the community survey after we had completed coding every interview, but before the first author had finished auditing coding decisions. Each of the first three authors independently assembled a list of survey questions, which the first author assembled into a single survey in Qualtrics (Qualtrics, 2024). Excluding consent and eligibility, the survey had 85 yes/no, multiple-choice, and Likert scale questions, which were grouped into 26 sections. To avoid priming participants to select certain responses, we randomly shuffled the order of multiple choice and yes/no options and randomly reversed the order of Likert scale items. Similar to our screening survey, certain questions were gated based on prior responses. We provide a complete copy of our survey, including its branching logic, in our Appendix (McCormack et al., 2026).
@@
Questions were either descriptive or exploratory. Figure 0 shows an example of a descriptive question. Each response option corresponds to the codes shown in Figure 0 for the theme “Why unsafe?”. Exploratory questions arose after we had finished conducting interviews, and they did not have a direct correspondence to the themes that we identified during qualitative analysis. For example, Höltervennhoff et al. (Höltervennhoff et al., 2023)’s study was published concurrent to our investigation, and three of their participants mentioned using either cargo-audit or cargo-deny to detect vulnerabilities in their dependencies. None of our interview participants had mentioned using these tools, but we included a survey question asking if respondents had used them before. We also relaxed our eligibility criteria for the survey to include individuals who would have been eligible for Höltervenhoff et al. (Höltervennhoff et al., 2023)’s study. Candidates were eligible to participate in their study if they had committed unsafe code to GitHub, but there was no indication if they had done so regularly. Our interview participants needed to have “regularly” engaged with unsafe code, but our survey respondents were eligible if they had “written, edited, read, audited, or engaged with unsafe Rust code in any way.”
@@
Two of our interview participants piloted an initial draft of the survey, and we incorporated their feedback into the final version. We distributed our community survey on September 22nd, 2023 to the same acquaintances and communities where we distributed our screening survey. We also published a link to our survey in the September 28th edition of the “This Week in Rust” newsletter and we advertised the survey during a presentation and poster sessions at an industry partners conference hosted by our institution. At the end of the survey, participants could provide a link to their profile on either GitHub or the Rust Programming Language Forum to have a chance to receive one of two \$250 gift cards. We randomly selected two participants who had provided a link to their profile with account activity prior to the start date of the survey.
We configured Qualtrics (Qualtrics, 2024) to reject multiple responses from the same individual. We also enabled reCaptcha (Liu, 2018) and RelevantID (RelevantID, 2026) to detect possibly fraudulent submissions. RelevantID uses several forms of metadata, including location and IP information; none of this information was retained or visible to the investigators. In our analysis, we excluded responses that had a reCaptcha score of less than 0.5, indicating that the response was likely from a bot. For RelevantID, we excluded responses that had a fraud score greater than 0.3 or were marked as likely to be duplicate.
@@ 4. Ethics & Threats to Validity
All procedures in this study were approved by our institution’s IRB. Before each intervention, participants were read or presented a script outlining the procedures of the study and they were asked to affirm consent. Participants were reminded to avoid sharing personally identifying information in their responses, and interview participants were asked to remain in a private space. The first author reviewed the contents of the replication package and redacted all direct (e.g. name, affiliation) and indirect (e.g. position, project) personal identifiers.
@@
Recruitment via snowball sampling and through online communities may bias our results toward certain areas of the Rust community. We attempted to maintain a diverse survey and interview population by sampling from multiple forums. However, both Reddit and “This Week in Rust” were overrepresented compared to the other communities where we distributed our survey. Our sample does not include developers who are not active in these communities. We cannot determine if our interview population has a similar bias, since we did not record where these participants had first learned about our study.
The "silent" developers excluded from these populations may have different experiences and perspectives that are not represented in the current findings. Rust’s unsafe “superpowers” are typically advertised as an advanced feature that developers “might run into every once in a while, but may not use every day” (Rust Community, 2023a). Developers who have engaged with unsafe code may more likely to have had more experience using Rust in any capacity than those who have not used unsafe code. We suspect that committed Rust experts are overrepresented in our interview population, since many were affiliated with prominent open source projects or had experience using the language within security-critical applications.
It is more difficult to evaluate the capabilities of our survey respondents. There was significant variance in the reported number of years of experience. However, on average, our survey population had nearly the same number of years of experience using Rust as our interview population, and more than a decade of experience within the field of software engineering. Most respondents reported behaviors that indicate non-trivial experience with Rust, such as using bug-finding or auditing tools and committing unsafe code to a published library. For this reason, we expect that our current results reflect the “best-case scenario” for developers’ experiences using unsafe Rust. Any aspects of Rust development that our population found to be difficult seem likely to be even more challenging for the broader Rust community.
Additionally, the gift cards that we offered for compensation are most easily redeemable in North America, which may have dissuaded certain individuals from participating in either stage of our study. We did not collect location data from participants, so we cannot determine if our results are biased toward individuals from any particular area.
@@
Respondents were eligible to complete our survey if they had one year of experience and had “engaged” with unsafe code in any way. A subset may have been eligible without having significant practical experience with unsafe code. We mitigated this by gating questions based on whether participants had used certain features. When relevant, we compare results from the entire population with results from the subset who “regularly” wrote or edited unsafe code, which matches the eligibility criteria for our interviews. This criteria may have introduced self-selection bias among our interview participants, since individuals who frequently use unsafe code may be more likely to perceive that it is necessary for their purposes. We did not enable reCaptcha (Liu, 2018), RelevantID (RelevantID, 2026), or any of Qualtrics’ other safeguards when recruiting participants for interviews. These safeguards are subject to false positives (Zhang et al., 2022), and may have flagged legitimate participants.
The validity of our results depends on participants’ ability to accurately self-report their behavior, and what developers perceive about their use of unsafe code may differ from how it is used in practice. We did not evaluate whether participants were able to reason correctly about unsafe code; only how they perceived their usage of these features. Höltervennhoff et al. (Höltervennhoff et al., 2023) found that developers who use unsafe code have varying degrees of awareness about its implications for correctness; 10 of their interview participants indicated that unsafe contexts entirely disable Rust’s static safety restrictions, which is incorrect. Our qualitative results are also unavoidably influenced by the individual biases of each investigator. We tried to mitigate this by auditing coding decisions and meeting frequently to resolve disagreements. Although most of our interview protocol was written to be neutral toward unsafe code, certain questions may have biased participants toward providing responses that favored safe coding practices, such as asking participants to describe how they “reason about memory safety” and “navigate the differences” between Rust and other languages.
@@ 5. Results
We report the demographics of interview and survey participants in Section 0. Then, we describe how these developers reason about foreign functions in Section 0 (RQ1), and we discuss the tooling they use in Section 0 (RQ2). We describe developers’ motivations for using unsafe code in Section 0 (RQ3), and we report how they reason about encapsulation in Section 0 (RQ4). We summarize key results at the end of each section.
@@ 5.1. Demographics
Table 1. Demographics of interview and survey participants.
(a) Interview participants’ affiliations ("Affil.") and years of experience using Rust ("Years") by ID.
ID Affil. Years
1 I 2
2 I,R 10
3 I,O 5
4 I,O 8
5 I,O 2
6 A 3
7 I 6
8 I 5
9 O 2
10 O 2
ID Affil. Years
11 I 5
12 I,A 6
13 I 10
14 R,O,I 7
15 I 1.5
16 I 3
17 A 4
18 I 3
19 I 4
I = Industry   R = Rust Team
A = Academia   O = Open Source
(b) The mean ($\mu$) and standard deviation ($\sigma$) of survey respondents’ years of experience in the field of software engineering ("SE"), as well as using Rust, C and C++. Entries are grouped by where respondents indicated that they had first heard about the survey. Respondents could select multiple options.
SE C C++ Rust
Location # $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
Rust Forums 16 17.3 10.1 7.5 6.8 5.6 6.2 4.5 2.4
Rust Discord 10 8.0 3.7 2.7 2.8 3.4 2.3 3.6 2.0
Reddit (/r/rust) 83 13.9 9.1 7.6 7.2 6.0 6.3 4.1 2.2
"This Week in Rust" 52 15.1 9.7 6.6 7.7 5.7 6.9 4.1 2.3
Other 3 23.7 15.8 17.3 14.6 9.0 5.6 3.7 1.5
All 160 14.5 9.5 7.2 7.5 5.8 6.3 4.1 2.2
@@ 19 I 4
Table 0 provides summary statistics on interview and survey participants. We invited all 42 eligible candidates to participate in interviews, and we interviewed 19 who responded to our invitation. Table 0 summarizes interview participants’ demographic information. There were 15 who had experience in industry, six in open source development, three in academia, and two were members of the Rust Team. They had 4.7 years of experience with Rust on average, and the most experienced participants had been using Rust for 10 years. We did not collect any additional demographic information from our interview participants.
@@
We received 368 survey responses, of which 240 were complete, 203 met our eligibility criteria, and 160 passed each of our measures for fraud detection. Hereafter, when we refer to “respondents,” we include only the 160 that passed each of our checks. We collected several categories of demographic information from survey respondents. The majority were young, educated, identified as male, and had experience in either industry or open source development. Nearly 66% had completed at least a bachelor’s degree, and 30% had completed a graduate degree. Most of the respondents (83%) were under 40 years of age, and 56% were between the ages of 18 and 29. We followed Spiel et al.’s (Spiel et al., 2019) guidelines for ethically surveying respondents’ gender; 82% identified as “Man”, 6% identified as “Non-binary”, 3% identified as “Woman” and 9% did not disclose their gender. Table 0 summarizes respondents’ years of experience with the field of software engineering, Rust, and other systems languages. On average, respondents had more than a decade of experience in software engineering, more than 5 years of experience with C and C++, and just over 4 years of experience using Rust.
@@
Table 2. How frequently survey respondents “engaged with” and wrote unsafe code. Percentages on the left include “Less than once a year” and “Yearly”, while those on the right include “Weekly” and “Daily”. The number of respondents in the sample is listed in parentheses on the righthand side.
ID Question Sample Distribution
2a How frequently do you engage with unsafe Rust code in any way? All 23%41%(160)
@@
2b How frequently do you write new Rust code or edit existing Rust code within an unsafe block or function? 31%26%(160)
Less than once a year    Yearly    Monthly    Weekly    Daily
@@
Participants had a variety of roles within the Rust community and differing levels of engagement with unsafe code. Among survey respondents, 74% had committed unsafe code to a public GitHub repository, and 66% had contributed to a published crate. Our interview population had regularly written or edited unsafe code, but 47% of our survey population did not. Since “regularly” is subjective, we asked the questions shown in Table 0 to provide a more specific time-frame and to distinguish editing unsafe code from practices that do not require code contributions, such as auditing. Only 26% (2) of survey respondents would write or edit unsafe code more than once a month, but 41% (2) would engage with it at this frequency.